OpenStudy (anonymous):

Use the quadratic formula to solve this equation 5j^2-20=0

OpenStudy (anonymous):

\[j = \pm 2\]

OpenStudy (anonymous):

5(j^2 - 4) = 0 j^2 -4 = 0 j^2 = 4 j = + - 2.

OpenStudy (anonymous):

no need for the formula j^2 = 4

OpenStudy (anonymous):

5j^2-20=0 5j^2=20 j^2=4